We are in a world of no corporate taxes. Markets in finance and investments are efficient. The risk-free rate of interest is 2% and the expected equity premium is 5%. In the competitive market for Waste Disposal Services, all company operate extremely efficiently. One such company is All Clean Disposals (ACD). Their gearing ratio is currently 50% and you can assume that their debt is riskless. Each year, in perpetuity. the firm generates operating income with the following probabilities. £100,000 with probability of 1/4, £250,000 with probability of 1/2 and £400,000 with probability of 1/4. The firm overall has an expected return of 6.25% What is the total value of the firm ACD? What is the expected return of ACDs equity? (10 marks) (b) Suppose ACD issues £750,000 of Equity and repurchases the same amount of Debt What is the operating income payable to shareholders now in each state? What is the overall return to ACDs equity now? Explain your answer (15 marks) (c) The Modigliani-Miller theory of capital structure is based on highly restrictive assumptions and hence has little practical use. (25 marks) Total 50 marks To what extent do you agree with this statement?
Answer #1

a ) Value of Operating Income every year = =(0.25*100000+0.5*250000+0.25*400000) = $ 250000.

Total Value of ACD =[(250000/ 6.5%)-debt] = $ 3,846,153.8- debt.

Expected Level of ACDs equity = Risk Free rate + Equity Premium = 2% +5% =7%.

b ) If ACD issues $ 750,000 worth of Equity and and repurchases debt of same value.

The interest payable on $ 750000 debt would be saved,

Gearing Ratio 50% (That means equity is double of debt)
Expected return on Equity 7
Total Expected Rate 6.5
6.5=(0.33X+0.67*7)
6.5 = (0.33X+ 4.69)

0.33X=1.89

X=1.89/0.33 = 5.73 %

Expected return on debt = 5.73%

Interest saved = 750000*5.73% = $ 42975.

Amount to be paid to equity holders now = 250000+42975 =292975 (Total Debt was not provided to deduct interest payable for debt).

C ) Yes, the MM Theory is based on certiain assumptions like.

  • No taxes
  • No transaction costs
  • No bankruptcy costs
  • Equivalence in borrowing costs for both companies and investors
  • Symmetry of market information, meaning companies and investors have the same information
  • No effect of debt on a company's earnings before interest and taxes

In the practical world, hardly any of the key assumptions hold true, in the real world there are taxes which should be factored in before doing any transactions,

Transaction Costs are prevalent, Bankruptcy has its own costs.

Cost of borrowing will defer a lot for companies and retail investors.

Information is not as symmetric as we would like to be.....etc